# Egorov-Type Semiclassical Limits for Open Quantum Systems with a Bi-Lindblad Structure

### v1

@@ Abstract
This paper develops a bridge between bi-Hamiltonian structures of Poisson–Lie type, contact Hamiltonian dynamics, and the Gorini–Kossakowski–Sudarshan–Lindblad (GKSL) formalism for quantum open systems. On the classical side, we consider bi-Hamiltonian systems defined by a Poisson pencil with non-trivial invariants. Using an exact symplectic realization, these invariants are lifted and projected onto a contact manifold, yielding a completely integrable contact Hamiltonian system in terms of dissipated quantities and a Jacobi-commutative algebra of observables. On the quantum side, we introduce a class of contact-compatible Lindblad generators: GKSL evolutions whose dissipative part preserves a commutative $C^{\ast}$-subalgebra generated by the quantizations of the classical dissipated quantities, and whose Hamiltonian part admits an Egorov-type semiclassical limit to the contact dynamics. This construction provides a mathematical mechanism compatible with the semiclassical limit for pure dephasing, compatible with integrability and contact dissipation. An explicit Euler-top-type Poisson–Lie pencil, inspired by deformed Euler top models, is developed as a fully worked-out example illustrating the resulting bi-Lindblad structure and its semiclassical behavior.
@@ 1 Introduction
Compatibility ensures the existence of a recursion operator (on a symplectic leaf where one of the Poisson structures is non-degenerate) and hierarchies of commuting Hamiltonians, which in turn give rise to rich algebraic and geometric structures and, under suitable regularity assumptions, to complete integrability in the sense of Liouville; see, for example, [M.M1984, C.M.P1993, M.C.F+1997, F.M.P2000, Fernandes1994, G.M.S2003].
A natural and geometrically significant source of bi-Hamiltonian systems is provided by Poisson–Lie groups. A Poisson tensor $\pi$ on a Lie group $G$ is called multiplicative if the group multiplication $m\colon G\times G\longrightarrow G$ is a Poisson map from $(G\times G,\pi\oplus\pi)$ to $(G,\pi)$. Multiplicative Poisson structures and their infinitesimal counterparts, Lie bialgebras, were introduced by Drinfel’d in the context of quantum groups [Drinfeld1988, Drinfeld1990] and further developed in the foundational work of Semenov–Tian–Shansky on the classical $r$-matrix formalism, Poisson pencils and factorization dynamics [STS1983, STS1985]. The global geometry of Poisson–Lie groups, including dressing transformations and symplectic groupoids, was elucidated in [LuWeinstein1990].
@@
then $X$ is a bi-Hamiltonian vector field whose integrability properties reflect the algebraic structure of the underlying Lie bialgebra and classical $r$-matrix. This mechanism underlies many classical integrable models on Lie groups and Poisson–Lie groups [STS1983, STS1985, KS2004, KS2008, LuWeinstein1990].
On the other hand, contact geometry provides a natural framework for describing non-Hamiltonian classical dynamics in a plethora of contexts, including certain dissipative systems and thermodynamically relevant flows; see, e.g., [d.L2019a, d.L2019b, G.G2022a, G.G2024a, Grabowski2013a, Lopez-Gordon2024]. There is a one-to-one correspondence between contact manifolds and homogeneous symplectic manifolds, which has been known for a while [Arnold1978, L.M1987], and whose applications to contact Hamiltonian dynamics have been recently analyzed in detail by Grabowska and Grabowski [Grabowski2013a, G.G2022a, G.G2023, G.G2024a]. We have exploited this correspondence in order to establish a novel characterisation of completely integrable contact Hamiltonian systems as completely integrable Hamiltonian systems (in the standard Liouville sense) that satisfy certain homogeneity conditions [C.d.L+2023, Lopez-Gordon2024, Colombo2025].
@@
In particular, in [Colombo2025], it is shown, in concrete examples, that one may take a Poisson structure on the symplectization compatible with the canonical one, use the corresponding recursion operator to construct a maximal family of homogeneous functions in involution, and then “project” these functions to a $(2n+1)$-dimensional contact manifold as functions in Jacobi-involution. What turns out to be too rigid for contact integrability is not Poisson compatibility on the symplectization, but the use of compatible Jacobi structures (see [NunesdaCosta1998a, M.M.P1999, C.D.N2010, P.N2003, Petalidou2002]) directly on the original contact manifold: compatible Jacobi pairs do not yield a recursion operator with enough independent eigenvalues to produce a full set of contact integrals.
@@
  • Systems defined on Poisson–Lie groups, and on any of their symplectic or exact symplectic realizations, are entirely classical Hamiltonian systems. Their integrability is governed by Poisson, Poisson–Nijenhuis and bi-Hamiltonian structures in the usual sense.
@@
where $\{\cdot,\cdot\}_{\alpha}$ is the Jacobi bracket defined by $\alpha$. The trivial symplectization of $(C,\alpha,h)$ is given by $(M=C\times(0,+\infty),\omega=\mathrm{d}(r\alpha),H=rh)$, where $r$ denotes the global coordinate of $(0,+\infty)$. There is a Lie algebra isomorphism between the Jacobi bracket $\{\cdot,\cdot\}_{\alpha}$ and the Poisson bracket $\{\cdot,\cdot\}_{\omega}$, which implies that conserved quantities for $(C,\alpha,h)$ are conserved quantities for $(M,\omega,H)$.
A symplectic form $\omega$ on a manifold $M$ is exact (i.e, there exists a one-form $\theta$ such that $\mathrm{d}\theta=-\omega$) if and only if there exists a Liouville vector field $\Delta$ on $M$ such that $\mathcal{L}_{\Delta}\omega=\omega$. Indeed, one can define $\theta=-\iota_{\Delta}\omega$. A tensor field $A$ on $M$ is called homogeneous of degree $w\in\mathbb{R}$ if $\mathcal{L}_{\Delta}A=w\cdot A$. On each hypersurface $C$ of $M$ which is transverse to $\Delta$, the restriction $\alpha={\left.\kern-1.2pt\theta\right|_{C}}$ is a contact form, and homogeneous functions of degree $1$ on $M$ project onto functions on $C$. Conversely, given a function $g$ on $C$, one can extend it, via the flow $\phi_{t}$ of $\Delta$, to a homogeneous function $\hat{g}$ on $M$ given by $\hat{g}(\phi_{t}(x))=e^{t}g(x)$ for each $x\in C$.
By quantum dissipation we refer to the irreversible, non-unitary component of the reduced dynamics of an open quantum system, described at the Markovian level by a GKSL (Lindblad) generator. This dissipative term arises from tracing out environmental degrees of freedom and accounts for phenomena such as decoherence, relaxation and thermalisation [BreuerPetruccione]. While such dissipation has no direct analog in closed Hamiltonian mechanics, it admits effective classical limits. In the present work we exhibit a class of GKSL generators whose semiclassical limit is naturally described by contact Hamiltonian dynamics, with dissipated quantities playing the role of robust invariants of the resulting reduced evolution.
@@
Before introducing the specific quantization scheme, it is useful to explain why such a quantum layer is natural in the present setting. A bi-Hamiltonian Poisson–Lie system produces, through an exact and homogeneous symplectic realization, a completely integrable contact system whose dynamics is intrinsically non-Hamiltonian, as contact vector fields does not preserve either the contact form or the associated “Hamiltonian” function. From the viewpoint of physics, contact systems appear precisely as effective or reduced descriptions of systems interacting with an environment, where certain degrees of freedom are not explicitly modelled. Quantum open systems [BreuerPetruccione, Lindblad1976], governed by Lindblad master (GKSL) equations, arise from a conceptually analogous mechanism: a unitary evolution on a larger Hilbert space is reduced to a non-unitary dynamics on the subsystem of interest.
@@
The classical part of our construction proceeds as follows. Starting from a bi-Hamiltonian Poisson–Lie system, we consider an exact symplectic realization where the Hamiltonian is homogeneous of degree one with respect to the Liouville vector field. On this exact symplectic manifold we exploit bi-Hamiltonian techniques to build a family of commuting integrals. Homogeneity then allows us to project the dynamics and its integrals to a contact manifold, obtaining a completely integrable contact system in the sense of [C.d.L+2023, Lopez-Gordon2024].
@@
Our answer is affirmative. We introduce the notion of a contact-compatible Lindblad generator: a GKSL dynamics whose Hamiltonian and dissipative parts are organized so that
@@
We also introduce a notion of bi-Lindblad structure, inspired by bi-Hamiltonian geometry: a pair of Lindblad generators whose convex combinations remain Lindbladian and whose common invariant observables encode the classical integrals of a bi-Hamiltonian system on a Poisson–Lie group. This provides a geometric bridge between Poisson–Lie geometry, contact integrable dynamics and physically admissible quantum dissipation.
An Euler-top-type Poisson–Lie system, inspired by the deformed Euler top introduced in [ballesteros2017poisson, G.I.M+2023], is developed here as a fully explicit example illustrating the classical–quantum correspondence between bi-Hamiltonian contact dynamics and Lindblad open-system evolution.
@@ 2 Bi-Hamiltonian geometry and contact integrability
This section reviews classical bi-Hamiltonian geometry and homogeneous Hamiltonian systems on exact symplectic manifolds, and recalls the notion of contact integrability introduced in [C.d.L+2023, Lopez-Gordon2024].
@@
On the one hand, bi-Hamiltonian geometry is entirely formulated in the Poisson (or symplectic) category: compatibility of Poisson structures leads to recursion operators and hierarchies of commuting Hamiltonians [M.M1984, M.C.F+1997, F.M.P2000, Fernandes1994, C.M.P1993, G.M.S2003]. On the other hand, homogeneity is a property with respect to the Liouville vector field on an exact symplectic manifold. In our context these two ingredients are logically independent: bi-Hamiltonian (or Poisson–Nijenhuis) structures will be used to generate integrals in involution in the usual symplectic sense, while homogeneity will be the geometric mechanism that allows us to project these integrals to a $(2n+1)$-dimensional contact manifold via symplectization.
@@ 2.1 Bi-Hamiltonian structures
On any symplectic leaf where $\Pi_{0}$ is nondegenerate, one can introduce the recursion operator (see, for instance, [M.M1984, M.C.F+1997, F.M.P2000, Fernandes1994])
@@
When the Nijenhuis torsion of $N$ vanishes, spectral invariants of $N$ yield hierarchies of commuting Hamiltonians, providing the standard bi-Hamiltonian framework for Liouville integrability [C.M.P1993, G.M.S2003]. This picture can be recast in the more general language of Poisson–Nijenhuis manifolds, Jacobi–Nijenhuis manifolds and their reductions [K.M1990, Wade2001, NunesdaCosta1998a, M.M.P1999, P.N2003, C.D.N2010, D.M.P2011], but in this section we restrict ourselves to the basic Poisson setting.
@@ 2.2 Exact symplectic manifolds and homogeneity
Throughout the paper we adopt the convention that an exact symplectic manifold $(M,\theta)$ carries the symplectic form $\omega\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=-\mathrm{d}\theta$. Hamiltonian vector fields are defined by $\iota_{X_{H}}\omega=\mathrm{d}H$, and the Liouville vector field $\Delta$ is characterized by $\iota_{\Delta}\omega=-\theta$.
@@
The Liouville vector field $\Delta$ is defined by $\iota_{\Delta}\omega=-\theta$.
@@
Following [C.d.L+2023, Lopez-Gordon2024], we say that a homogeneous Hamiltonian system is integrable if it admits a full set of independent, homogeneous first integrals in involution.
@@
In our later construction, we will start from a Poisson (or Poisson–Lie) bi-Hamiltonian system and, when an exact symplectic realization is available, we will require the Hamiltonian and its integrals to be $1$-homogeneous with respect to the corresponding Liouville vector field. Compatibility of Poisson tensors is responsible for integrability in the Liouville sense, whereas homogeneity is what will allow us to relate this symplectic picture with a contact Hamiltonian system on a $(2n+1)$-dimensional manifold via symplectization. These are therefore two independent geometric requirements that must be arranged simultaneously in our applications.
@@ 2.3 Contact Hamiltonian systems and integrability
Given a function $h\in\mathscr{C}^{\infty}(C)$, the associated contact Hamiltonian vector field $X_{h}$ is defined by (see, e.g., [d.L2019b, G.G2022a, Lopez-Gordon2024])
@@
The contact form $\alpha$ induces a Jacobi structure $(\Lambda,R)$ on $C$ [Lichnerowicz1977a, Lichnerowicz1978, d.L.M+2003, G.U2003], and hence a Jacobi bracket on $\mathscr{C}^{\infty}(C)$ given by
@@
This notion was introduced in [G.G.M+2020a] and further developed in [d.L2019a, C.d.L+2023, Lopez-Gordon2024] as the natural replacement, in contact geometry, of conserved quantities in Hamiltonian dynamics.
@@
In this language, the contact Liouville–Arnol’d theorem proved in [C.d.L+2023] extends the classical symplectic theory [Arnold1978, Audin2004] to the contact category. In particular, completely integrable contact systems admit local action–angle type coordinates adapted to the Reeb dynamics and to the dissipated quantities.

### v2

@@ Abstract
This paper develops a bridge between bi-Hamiltonian structures of Poisson–Lie type, contact Hamiltonian dynamics, and the Gorini–Kossakowski–Sudarshan–Lindblad (GKSL) formalism for open quantum systems. On the classical side, we consider bi-Hamiltonian systems defined by a Poisson pencil with non-trivial invariants. Using an exact symplectic realization, these invariants are lifted and projected onto a contact manifold, yielding a completely integrable contact Hamiltonian system and a Jacobi-commutative algebra of observables. On the quantum side, we introduce a class of contact-compatible Lindblad generators: GKSL evolutions whose dissipative part preserves a commutative $C^{\ast}$-subalgebra generated by the quantizations of the classical dissipated quantities, and whose Hamiltonian part admits an Egorov-type semiclassical limit to the contact dynamics. This construction provides a mathematical mechanism compatible with the semiclassical limit for pure dephasing, compatible with integrability and contact dissipation. An explicit Poisson–Lie pencil, inspired by deformed Euler top models, is developed as a fully worked-out example illustrating the resulting bi-Lindblad structure and its semiclassical behavior.
@@ 1 Introduction
Compatibility ensures the existence of a recursion operator (on a symplectic leaf where one of the Poisson structures is non-degenerate) and hierarchies of commuting Hamiltonians, which in turn give rise to rich algebraic and geometric structures and, under suitable regularity assumptions, to complete integrability in the sense of Liouville; see, for example, [46, 8, 45, 21, 22, 26].
A natural and geometrically significant source of bi-Hamiltonian systems is provided by Poisson–Lie groups. A Poisson tensor $\pi$ on a Lie group $G$ is called multiplicative if the group multiplication $m\colon G\times G\longrightarrow G$ is a Poisson map from $(G\times G,\pi\oplus\pi)$ to $(G,\pi)$. Multiplicative Poisson structures and their infinitesimal counterparts, Lie bialgebras, were introduced by Drinfel’d in the context of quantum groups [19, 20] and further developed in the foundational work of Semenov–Tian–Shansky on the classical $r$-matrix formalism, Poisson pencils and factorization dynamics [57, 56]. The global geometry of Poisson–Lie groups, including dressing transformations and symplectic groupoids, was elucidated in [44].
@@
then $X$ is a bi-Hamiltonian vector field whose integrability properties reflect the algebraic structure of the underlying Lie bialgebra and classical $r$-matrix. This mechanism underlies many classical integrable models on Lie groups and Poisson–Lie groups [57, 56, 37, 38, 44].
On the other hand, contact geometry provides a natural framework for describing non-Hamiltonian classical dynamics in a plethora of contexts, including certain dissipative systems and thermodynamically relevant flows; see, e.g., [14, 15, 27, 29, 43]. There is a one-to-one correspondence between contact manifolds and homogeneous symplectic manifolds, which has been known for a while [1, 39], and whose applications to contact Hamiltonian dynamics have been recently analyzed in detail by Grabowska and Grabowski [31, 27, 28, 29]. We have exploited this correspondence in order to establish a novel characterisation of completely integrable contact Hamiltonian systems as completely integrable Hamiltonian systems (in the standard Liouville sense) that satisfy certain homogeneity conditions [10, 43, 11].
At this stage, it is natural to ask why one should restrict to homogeneous symplectic realizations projectable to contact manifolds, rather than considering a generic symplectic realization of an open quantum system. Although a classical limit can be obtained from such generic symplectic realizations, these constructions are highly non-unique and do not provide a canonical description of dissipation. Requiring the realization to be homogeneous and projectable to a contact structure selects a distinguished geometric class in which dissipation, contraction of phase-space volume, and irreversibility are encoded intrinsically. In this framework, the effective non-unitary dynamics arises as the projection of a unitary flow on a higher-dimensional homogeneous symplectic manifold, yielding a robust and universal semiclassical limit. The contact-compatible structure is therefore not an additional assumption, but the minimal geometric setting capturing the correct physics of open quantum systems.
@@
In particular, in [11], it is shown, in concrete examples, that one may take a Poisson structure on the symplectization compatible with the canonical one, use the corresponding recursion operator to construct a maximal family of homogeneous functions in involution, and then “project” these functions to a $(2n+1)$-dimensional contact manifold as functions in Jacobi-involution. What turns out to be too rigid for contact integrability is not Poisson compatibility on the symplectization, but the use of compatible Jacobi structures (see [53, 47, 52, 55]) directly on the original contact manifold: compatible Jacobi pairs do not yield a recursion operator with enough independent eigenvalues to produce a full set of contact integrals.
@@
  • Systems defined on a Poisson–Lie group $G$, and on any of its symplectic or exact symplectic realizations, are entirely classical Hamiltonian systems. Their integrability is governed by Poisson, Poisson–Nijenhuis and bi-Hamiltonian structures in the usual sense.
@@
where $\{\cdot,\cdot\}_{\alpha}$ is the Jacobi bracket defined by $\alpha$. The trivial symplectization of $(C,\alpha,h)$ is given by $(M=C\times(0,+\infty),\omega=\mathrm{d}(r\alpha),H=rh)$, where $r$ denotes the global coordinate of $(0,+\infty)$. There is a Lie algebra isomorphism between the Jacobi bracket $\{\cdot,\cdot\}_{\alpha}$ and the Poisson bracket $\{\cdot,\cdot\}_{\omega}$, which implies that dissipated quantities for $(C,\alpha,h)$ are in one-to-one correspondence with conserved quantities for $(M,\omega,H)$.
A symplectic form $\omega$ on a manifold $M$ is exact (i.e., there exists a one-form $\theta$ such that $\mathrm{d}\theta=-\omega$) if and only if there exists a Liouville vector field $\Delta$ on $M$ such that $\mathcal{L}_{\Delta}\omega=\omega$. Indeed, one can define $\theta=-\iota_{\Delta}\omega$. A tensor field $A$ on $M$ is called homogeneous of degree $w\in\mathbb{R}$ if $\mathcal{L}_{\Delta}A=w\cdot A$. On each hypersurface $C$ of $M$ which is transverse to $\Delta$, the restriction $\alpha={\left.\kern-1.2pt\theta\right|_{C}}$ is a contact form, and homogeneous functions of degree $1$ on $M$ project onto functions on $C$. Conversely, given a function $g$ on $C$, one can extend it, via the flow $\phi_{t}$ of $\Delta$, to a homogeneous function $\hat{g}$ on $M$ given by $\hat{g}(\phi_{t}(x))=e^{t}g(x)$ for each $x\in C$.
By quantum dissipation we refer to the irreversible, non-unitary component of the reduced dynamics of an open quantum system, described at the Markovian level by a GKSL (Lindblad) generator. This dissipative term arises from tracing out environmental degrees of freedom and accounts for phenomena such as decoherence, relaxation and thermalisation [6]. While such dissipation has no direct analog in classical Hamiltonian mechanics, it admits effective classical limits. In the present work we exhibit a class of GKSL generators whose semiclassical limit is naturally described by contact Hamiltonian dynamics, with dissipated quantities playing the role of robust invariants of the resulting reduced evolution.
@@
Before introducing the specific quantization scheme, it is useful to explain why such a quantum layer is natural in the present setting. A bi-Hamiltonian Poisson–Lie system produces, through an exact and homogeneous symplectic realization, a completely integrable contact system whose dynamics is intrinsically non-Hamiltonian, as contact vector fields does not preserve either the contact form or the associated “Hamiltonian” function. From the viewpoint of physics, contact systems appear precisely as effective or reduced descriptions of systems interacting with an environment, where certain degrees of freedom are not explicitly modelled. Quantum open systems [6, 42], governed by Lindblad master (GKSL) equations, arise from a conceptually analogous mechanism: a unitary evolution on a larger Hilbert space is reduced to a non-unitary dynamics on the subsystem of interest.
@@
The classical part of our construction proceeds as follows. Starting from a bi-Hamiltonian Poisson–Lie system, we consider an exact symplectic realization where the Hamiltonian is homogeneous of degree one with respect to the Liouville vector field. On this exact symplectic manifold we exploit bi-Hamiltonian techniques to build a family of commuting integrals. Homogeneity then allows us to project the dynamics and its integrals to a contact manifold, obtaining a completely integrable contact system in the sense of [10, 43].
@@
Our answer is affirmative (see Theorem 4.24). We introduce the notion of a contact-compatible Lindblad generator: a GKSL dynamics whose Hamiltonian and dissipative parts are organized so that
@@
We also introduce a notion of bi-Lindblad structure, inspired by bi-Hamiltonian geometry: a pair of Lindblad generators whose convex combinations remain Lindbladian and whose common invariant observables encode the classical integrals of a bi-Hamiltonian system on a Poisson–Lie group. This provides a geometric bridge between Poisson–Lie geometry, contact integrable dynamics and physically admissible quantum dissipation (see Proposition 4.23).
An Euler-top-type Poisson–Lie system, inspired by the deformed Euler top introduced in [3, 33], is developed here as a fully explicit example illustrating the classical–quantum correspondence between bi-Hamiltonian contact dynamics and Lindblad open-system evolution.
@@ 2 Bi-Hamiltonian geometry and contact integrability
This section reviews classical bi-Hamiltonian geometry and homogeneous Hamiltonian systems on exact symplectic manifolds, and recalls the notion of contact integrability introduced in [10, 43].
@@
On the one hand, bi-Hamiltonian geometry is entirely formulated in the Poisson (or symplectic) category: compatibility of Poisson structures leads to recursion operators and hierarchies of commuting Hamiltonians [46, 45, 21, 22, 8, 26]. On the other hand, homogeneity is a property with respect to the Liouville vector field on an exact symplectic manifold. These two ingredients are independent: bi-Hamiltonian (or Poisson–Nijenhuis) structures will be used to generate integrals in involution in the usual symplectic sense, while homogeneity will be the geometric mechanism that allows us to project these integrals to a contact manifold of one dimension less.
@@ 2.1 Bi-Hamiltonian structures
On any symplectic leaf where $\Pi_{0}$ is non-degenerate, one can introduce the recursion operator (see, for instance, [46, 45, 21, 22])
@@
The Nijenhuis torsion of $N$ vanishes, and its eigenvalues are commuting functions in involution providing the standard bi-Hamiltonian framework for Liouville integrability [8, 26].
@@ 2.2 Exact symplectic manifolds and homogeneity
Throughout the paper we adopt the convention that an exact symplectic manifold $(M,\theta)$ carries the symplectic form $\omega\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=-\mathrm{d}\theta$. Hamiltonian vector fields are defined by $\iota_{X_{H}}\omega=\mathrm{d}H$, and the Liouville vector field $\Delta$ is uniquely determined by $\iota_{\Delta}\omega=-\theta$.
@@
Following [10, 43], we say that a homogeneous Hamiltonian system is integrable if it admits a full set of independent, homogeneous first integrals in involution.
@@
In our later construction, we will start from a Poisson (or Poisson–Lie) bi-Hamiltonian system and, when an exact symplectic realization is available, we will require the Hamiltonian and its integrals to be $1$-homogeneous with respect to the corresponding Liouville vector field. Compatibility of Poisson tensors is responsible for integrability in the Liouville sense, whereas homogeneity is what will allow us to project this symplectic picture to a contact Hamiltonian system. These are therefore two independent geometric requirements that must be arranged simultaneously in our applications.
@@ 2.3 Contact Hamiltonian systems and integrability
Given a function $h\in\mathscr{C}^{\infty}(C)$, the associated contact Hamiltonian vector field $X_{h}$ is defined by (see, e.g., [15, 27, 43])
@@
The contact form $\alpha$ induces a Jacobi structure $(\Lambda,R)$ on $C$ [40, 41, 16], and hence a Jacobi bracket on $\mathscr{C}^{\infty}(C)$ given by
@@
This notion was introduced in [24] and further developed in [14, 10, 43, 25] as the natural replacement, in contact geometry, of conserved quantities in Hamiltonian dynamics.
@@
In this language, the contact Liouville–Arnol’d theorem proved in [10] extends the classical symplectic theory [1, 2] to the contact category. In particular, completely integrable contact systems admit local action–angle type coordinates adapted to the contact Hamiltonian dynamics and to a certain foliation defined by the dissipated quantities.